# Mathematical Foundations of Polyphonic Music Generation via Structural Inductive Bias A Monograph

### v2

@@ 9.6 Conclusion
In this paper, we identified the SVD Paradox and provided a mathematical resolution via the RPTP theory. We demonstrated that structural sparsity, when combined with transversal mixing, acts as a rank-expanding mechanism rather than a constraint. Our contribution is threefold:
@@
    Methodological: We introduced Smart v3 (Info-Theoretic), proving that data-driven topological design allows for extreme parameter compression (13%) without sacrificing representational capacity.
Our Uniqueness Theorem and experimental results suggest that the Shuffled Block-Diagonal architecture (specifically in its NMI-optimized form) is the Canonical Form for future large-scale deep learning, offering a theoretically grounded path to scaling beyond the density bottleneck.
@@ B.2 Justification of the Scaling Assumption ($B\propto\sqrt{N_{params}}$)
    Initialization Schemes: Standard initialization methods, such as He Initialization [93] and Xavier/Glorot Initialization [94], normalize weights such that the variance is $O(1/N)$, leading to a total squared norm proportional to NN.

### v3

@@ 9.9 Conclusion
In this paper, we identified the SVD Paradox and provided a mathematical resolution via the RPTP theory. We demonstrated that structural sparsity, when combined with transversal mixing, acts as a rank-expanding mechanism rather than a constraint.
Our contribution is threefold:
@@
    Methodological: We introduced Smart v3 (Info-Theoretic) and its scalable realization Smart v4 (Hybrid), proving that data-driven topological design allows for extreme parameter compression (13%) without sacrificing representational capacity.
Our Uniqueness Theorem and experimental results suggest that the Shuffled Block-Diagonal architecture (specifically in its NMI-optimized form or its approximated variants) is the Canonical Form for future large-scale deep learning, offering a theoretically grounded path to scaling beyond the density bottleneck.
@@ B.2 Justification of the Scaling Assumption ($B\propto\sqrt{N_{params}}$)
    Initialization Schemes: Standard initialization methods, such as He Initialization [37] and Xavier/Glorot Initialization [30], normalize weights such that the variance is $O(1/N)$, leading to a total squared norm proportional to $N$.